### v1

@@ Abstract
Context. Classical Cepheids (DCEPs) are crucial for calibrating the extragalactic distance ladder, ultimately enabling the determination of the Hubble constant ($H_{0}$) through the period-luminosity ($PL$) and period-Wesenheit ($PW$) relations they exhibit. Hence, it’s vital to understand how the $PL$ and $PW$ relations depend on metallicity. This is the purpose of the C-MetaLL survey, within which this work is situated. DCEPs are also very important tracers of the young populations placed along the Galactic disc.
@@
Results. The sample presented in this paper was complemented with that already published in the context of the C-MetaLL survey, resulting in a total of 292 pulsators whose spectra have been analysed in a homogeneous way. These data were used to study the abundance gradients of the Galactic disc in a range of Galactocentric radii ($R_{GC}$) spanning the range 5-20 kpc.
Conclusions. For most of the elements we found a clear negative gradient, with a slope of $-0.071\pm 0.003$ dex kpc-1 for [Fe/H] case. Through a qualitative fit with the Galactic spiral arms we shown how our farthest targets ( $R_{GC}$10 kpc) trace both the Outer and OSC arms. The homogeneity of the sample will be of pivotal importance for the study of the metallicity dependance of the DCEP $PL$ relations.
@@
giant planet formation – $\kappa$-mechanism – stability of gas spheres
@@ 1 Introduction
Classical Cepheids (DCEPs) play a fundamental role as standard candles in determining extragalactic distances due to the Leavitt Law (Leavitt & Pickering, 1912), a relationship between period and luminosity ($PL$). Calibrated using independent distances based on geometric methods such as trigonometric parallaxes, eclipsing binaries, and water masers, these relations serve as the first step in constructing the cosmic distance scale. This way, they calibrate secondary distance indicators, including Type Ia supernovae (SNIa), allowing us to measure the distances of distant galaxies in the steady Hubble flow. The calibration of this three-step procedure, often referred to as the cosmic distance ladder, enables us to reach the Hubble flow and measure the Hubble constant $H_{0}$, which connects the distance and the recession velocity of galaxies (e.g. Sandage & Tammann, 2006; Freedman et al., 2012; Riess et al., 2016, and references therein).
Table 1: Main properties of the 180 program DCEPs
@@
Star Gaia_source_id RA Dec Period Mode $G$ $G_{BP}$ $G_{RP}$ $\varpi$ $\sigma$ Source Notes
deg deg days mag mag mag mas mas
AA_Gem 3430067092837622272 06:06:34.95 +26:19:45.2 11.31285 DCEP_F 9.363 9.956 8.62 0.2749 0.0177 ESPaDOnS SOS,SOS
AD_Pup 5614312705966204288 07:48:03.85 -25:34:40.0 13.59738 DCEP_F 9.541 10.125 8.786 0.2331 0.0165 ESPaDOnS SOS,SOS
AP_Sgr 4066429066901946368 18:13:02.50 -23:07:02.3 5.05808 DCEP_F 6.922 7.46 6.219 1.1815 0.024 ESPaDOnS P21,DR3
AQ_Pup 5597379741549105280 07:58:22.09 -29:07:48.3 30.15959 DCEP_F 8.127 8.974 7.25 0.2751 0.0226 ESPaDOnS SOS,SOS
ASAS-SN_J065046.50-085808.7 3050543819559526912 06:50:46.50 -08:58:08.8 6.94564 DCEP_F 13.49 14.402 12.547 0.1198 0.0179 UVES,P110 SOS,SOS
@@
 [^1] Meaning of the columns: Star = literature name of the DCEP; Gaia_source_id= $Gaia$ identifier; RA, Dec = equatorial coordinates at J2000; Period and Mode = period and mode of pulsation: DCEP_F, DCEP_1O and DCEP_MULTI refer to fundamental, first overtone and multimode DCEPs, respectively; $G$, $G_{BP}$ and $G_{RP}$ are the magnitudes in the $Gaia$ bands; $\varpi$ and $\sigma$ are the $Gaia$ parallas and its error; Source show the instrument used to observe the star A portion is shown here for guidance regarding its form and content. The machine-readable version of the full table will be published at the CDS (Centre de Données astronomiques de Strasbourg, https://cds.u-strasbg.fr/).
@@
In recent years a still unresolved debate regards the well-known discrepancy between the values of $H_{0}$ obtained by the SH0ES [^2] Supernovae, HO, for the Equation of State of Dark energy project through the cosmic distance ladder ($H_{0}=$73.01$\pm$0.99 km s-1 Mpc-1, Riess et al., 2022), and the value computed by the Planck Cosmic Microwave Background (CMB) project based on the flat $\Lambda$ Cold Dark Matter ($\Lambda$CDM) model ($H_{0}=$67.4$\pm$0.5 km s-1 Mpc-1, Planck Collaboration et al., 2020). If no solution is found for this intricate $H_{0}$ tension, this could hint for a revision of the $\Lambda$CDM model. It is therefore critical to evaluate the discrepancy through the analysis of the residual systematics (see e.g. Dainotti et al., 2021; Freedman, 2021; Riess et al., 2021, and reference therein).
@@
Metallicity may play an important role in being one of the possible residual sources of uncertainty in the cosmic distance ladder, influencing the calibration of the DCEP $PL$ relations. Since the shape and width of the DCEP instability strip are predicted to be affected by metallicity variations (e.g. Caputo et al., 2000), a direct consequence should be seen when estimating the $PL$ relations coefficients (Marconi et al., 2005, 2010; De Somma et al., 2022, and references therein). To avoid systematic effects in the calibration of the extragalactic distance scale (e.g. Romaniello et al., 2008; Bono et al., 2010; Riess et al., 2016), one has to consider the metallicity dependence of the $PL$ relations and the reddening-free Wesenheit magnitudes (Madore, 1982), also for those cases where this effect might be small, such as when dealing with the near-infrared regime (NIR, see e.g. Fiorentino et al., 2013; Gieren et al., 2018).
@@
In the past years, the lack of accurate independent distances has hampered direct empirical evaluations of the metallicity dependence of $PL$ relations using Galactic DCEPs with precise [Fe/H] measurements based on high-resolution (HiRes hereafter) spectroscopy, but this scenario was completely changed with the advent of the Gaia mission (Gaia Collaboration et al., 2016). Accurate parallaxes were provided with data release 2 (DR2, Gaia Collaboration et al., 2018) and further improved with the early data release 3 (EDR3, Gaia Collaboration et al., 2021), in addition to the discovery of hundreds of new Galactic DCEPs (Clementini et al., 2019; Ripepi et al., 2019, 2022a). A substantial sample is thus possible to be considered when other surveys are taken into consideration, such as the OGLE Galactic Disk survey (Udalski et al., 2018) and the Zwicky Transient Facility (ZTF, Chen et al., 2020), making it possible to improve studies not only related to the cosmic distance scale but also for Galactic studies (e.g. Lemasle et al., 2022; Trentin et al., 2023, and references therein). Another constraint was the limited availability of HiRes measurements confined to the solar neighbourhood. This restriction resulted in a narrow range of [Fe/H] values centred on solar or slightly supersolar values, with a small dispersion of 0.2-0.3 dex (e.g. Genovali et al., 2014; Luck, 2018; Groenewegen, 2018; Ripepi et al., 2019) Consequently, this limitation hindered the ability to generate results with statistical significance.
@@
To overcome these problems, we started a project named C-MetaLL [^3] https://sites.google.com/inaf.it/c-metall/home (Cepheid - Metallicity in the Leavitt Law, see Ripepi et al., 2021, for a full description), with the primary goal to measure the chemical abundance of a sample of at least 300 Galactic DCEPs through HiRes spectroscopy, as well as provide homogeneous multi-band time-series photometry. The main aim was to enlarge the iron abundance range towards the metal-poor regime —that is, [Fe/H]<absent<-< -0.4 dex— where only a few Galactic stars had abundance measurements in the literature. In the first two papers of the series (Ripepi et al., 2021; Trentin et al., 2023) (R21 and T23 from now on, respectively), we published accurate abundances for more than 25 chemical species for a total of 114 DCEPs, while the first results on the photometric part were presented in Bhardwaj et al. (2024) . In particular, in T23, we obtained measures for 43 objects with [Fe/H]<absent<-< -0.4 dex, reaching abundances as low as --1.1 dex. This paper represents the sixth manuscript in the series and a direct follow-up of T23, presenting the spectroscopic analysis of 180 new stars based on a total of 331 high-resolution spectra and the study of the galactic gradient for 29 chemical elements.
@@
While the main scope of this work is to study the metallicity dependence of the DCEP $PL$ relations in the context of the extragalactic distance scale, nevertheless, the sample of DCEPs presented in this work, in conjunction with our previous results, amounts to a total of 294 DCEPs with homogeneous derivation of stellar parameters and abundances. In addition, as we shall see later in this article, our sample is evenly distributed in metallicity, spanning a range from +0.5 dex to --1 dex in [Fe/H] and 5 to 20 kpc in Galactocentric radius. The mentioned characteristics make our sample unique and therefore a perfect test for the study of the Galactic disc abundance gradient, in comparison with literature works which in general use heterogeneous spectroscopic samples or datasets heavily unbalanced towards the solar vicinity (Luck, 2018; Lemasle et al., 2018; Minniti et al., 2020; Ripepi et al., 2022b; da Silva et al., 2023; Matsunaga et al., 2023).
@@
The paper is organized as follows: in Sect. 2 we describe the DCEPs sample and their properties; in Sect. 3 we describe the analysis technique; in Sect. 4 and Sect. 5, we describe and discuss our results and in Sect. 6 we outline our conclusions.
@@ 2.1 Sample
The sample of DCEPs presented in this work was selected from the DCEP catalogue published in $Gaia$ data release 3 (DR3 Ripepi et al., 2023) to maximize objects with low metallicity and long periods. The list of the 180 targets is shown in Table 1, where we list the main characteristics of the stars, including their period and mode of pulsation. The distribution of the targets along the MW disc is displayed in Fig. 1. The distribution of the program stars ranges from about 5 to 18 kpc in $R_{GC}$ and covers all four Galactic quadrants, with a significant concentration in the third. In the figure, the stars are colour-coded according to their periods. The period distribution seems to show short-period DCEPs at larger Galactic radii, especially in the third quadrant. This is not a coincidence, but it is because there exists a period-age relation for DCEPs (see e.g. Bono et al., 2005) so that shorter periods indicate larger ages. It has been found that at larger Galactocentric radii the DCEPs are on average older (e.g. Skowron et al., 2019; De Somma et al., 2020, 2021). This explains the lack of long-period DCEPs at large radii in our sample. An additional feature displayed in Fig. 1 is the presence of three stars with polar angles between 135o and 225o, namely Gaia_DR2_4087335043492541696, V532_Sco and ASAS-SN_J165340.10-332041.7. Given the peculiar position along the Galactic disc, these pulsators may be type II Cepheids. An in-depth analysis of these objects is provided in Appendix A. In addition, a fourth star, OGLE CEP-GD-0069 which was originally classified as DCEP in Udalski et al. (2018) turned out to be a BL Her variable according to Ripepi et al. (2023). Also this star is discussed in the appendix. Given their uncertain nature, these four stars have been excluded from the following analysis.
Figure 1: Map along the Galactic disc of the program stars. The points are colour-coded according to the pulsator’s period. The position of the sun is shown with a yellow-black circle.
@@
Star HJD Texp SNR $T_{eff}$ $\log g$ $\xi$ $V_{broad}$ $RV$ Source
days s K dex km/s km/s km/s
AA_Gem 59269.76384 750 102 5656 $\pm$ 68 0.92 $\pm$ 0.05 3.0 $\pm$ 0.2 11.0 $\pm$ 1.0 0.9 $\pm$ 0.3 ESPaDOnS
AA_Gem 59630.79632 410 92 5877 $\pm$ 66 1.22 $\pm$ 0.05 3.3 $\pm$ 0.2 12.0 $\pm$ 1.0 0.8 $\pm$ 0.2 ESPaDOnS
AA_Gem 59838.14990 500 80 4887 $\pm$ 68 0.1 $\pm$ 0.05 2.8 $\pm$ 0.2 9.0 $\pm$ 1.0 24.6 $\pm$ 0.2 ESPaDOnS
AD_Pup 59269.81690 750 78 5128 $\pm$ 100 0.38 $\pm$ 0.05 2.7 $\pm$ 0.2 6.0 $\pm$ 1.0 74.0 $\pm$ 0.3 ESPaDOnS
AD_Pup 59630.82225 410 77 5983 $\pm$ 83 0.58 $\pm$ 0.05 3.4 $\pm$ 0.3 10.0 $\pm$ 1.0 68.3 $\pm$ 0.2 ESPaDOnS
 [^4] The different columns report: the name of the star, the Heliocentric Julian Day at which the spectrum has been collected, the exposure time, the signal-to-noise ratio, the effective temperature, logarithm of the gravity, microturbolent velocity, broadening velocity, radial velocity, and the source (e.g. the instrument used to collect the data) of the high-resolution spectroscopy. A portion is shown here for guidance regarding its form and content. The machine-readable version of the full table will be published at the CDS (Centre de Données astronomiques de Strasbourg, https://cds.u-strasbg.fr/).
@@
 [^5] The table is divided in three parts. The first and last columns report the name of the star and the source, respectively (see Table 2 and 1). The other columns report the estimated abundances (and relative errors) in solar terms for the chemical species analysed in this work. The machine-readable version of the full table will be published at the CDS (Centre de Données astronomiques de Strasbourg, https://cds.u-strasbg.fr/).

### v3

@@ Abstract
Context. Classical Cepheids (DCEPs) are crucial for calibrating the extragalactic distance ladder, ultimately enabling the determination of the Hubble constant through the period-luminosity ($PL$) and period-Wesenheit ($PW$) relations that they exhibit. Hence, it is vital to understand how the $PL$ and $PW$ relations depend on metallicity. This is the purpose of the C-MetaLL survey, within which this work is situated. The DCEPs are also very important tracers of the young populations placed along the Galactic disc.
@@
Results. The sample presented in this paper was complemented with that already published in the context of the C-MetaLL survey, resulting in a total of 292 pulsators whose spectra have been analysed in a homogeneous way. These data were used to study the abundance gradients of the Galactic disc in a range of Galactocentric radii ($R_{GC}$) spanning the range of 5-20 kpc.
Conclusions. For most of the elements, we have found a clear negative gradient, with a slope of $-0.064\pm 0.003$ dex kpc-1 for [Fe/H] case. Through a qualitative fit with the Galactic spiral arms, we show how our farthest targets ( $R_{GC}$10 kpc) trace both the Outer and Outer Scutum-Centaurus arms. The homogeneity of the sample will be of pivotal importance for the study of the metallicity dependence of the DCEP $PL$ relations.
@@
stars: abundances – stars: distances – stars: fundamental parameters – stars: variables: Cepheids – Galaxy: disc
@@ 1 Introduction
Classical Cepheids (DCEPs) play a fundamental role as standard candles in determining extragalactic distances due to the Leavitt law (Leavitt & Pickering, 1912), a relationship between period and luminosity ($PL$). Calibrated using independent distances based on geometric methods such as trigonometric parallaxes, eclipsing binaries, and water masers, these relations serve as the first step in constructing the cosmic distance scale. This way, they calibrate secondary distance indicators, including type Ia supernovae (SNe Ia), allowing us to measure the distances of distant galaxies in the steady Hubble flow. The calibration of this three-step procedure, often referred to as the cosmic distance ladder, enables us to reach the Hubble flow and measure the Hubble constant, $H_{0}$, which connects the distance and the recession velocity of galaxies (e.g. Sandage & Tammann, 2006; Freedman et al., 2012; Riess et al., 2016, and references therein).
Table 1: Main properties of the 180 programme DCEPs
@@
Star Gaia_source_id RA Dec Period Mode $G$ $G_{BP}$ $G_{RP}$ $\varpi$ $\sigma$ $D$ $\sigma$ Source Notes
deg deg days mag mag mag mas mas kpc kpc
AA Gem 3430067092837622272 06:06:34.95 +26:19:45.2 11.31285 DCEP_F 9.363 9.956 8.62 0.2749 0.0177 3.81 0.11 ESPaDOnS SOS,SOS
AD Pup 5614312705966204288 07:48:03.85 -25:34:40.0 13.59738 DCEP_F 9.541 10.125 8.786 0.2331 0.0165 0.82 0.04 ESPaDOnS P21,DR3
AP Sgr 4066429066901946368 18:13:02.50 -23:07:02.3 5.05808 DCEP_F 6.922 7.46 6.219 1.1815 0.024 2.85 0.12 ESPaDOnS SOS,SOS
AQ Pup 5597379741549105280 07:58:22.09 -29:07:48.3 30.15959 DCEP_F 8.127 8.974 7.25 0.2751 0.0226 3.98 0.14 ESPaDOnS SOS,SOS
ASAS-SN J065046.50-085808.7 3050543819559526912 06:50:46.50 -08:58:08.8 6.94564 DCEP_F 13.49 14.402 12.547 0.1198 0.0179 2.66 0.10 UVES,P110 SOS,SOS
@@
 [^1] Meaning of the columns: Star = literature name of the DCEP. Gaia_source_id= $Gaia$ DR3 identifier. RA, Dec = equatorial co-ordinates at J2000. Period and Mode = period and mode of pulsation; DCEP_F, DCEP_1O, and DCEP_MULTI refer to fundamental, first overtone, and multimode DCEPs, respectively. $G$, $G_{BP}$, and $G_{RP}$ are the magnitudes in the $Gaia$ bands. $\varpi$ and $\sigma$ are the $Gaia$ parallax and its error. $D$ and $\sigma$ are the distances and its error. Source shows the instrument used to observe the star. Notes report the origin of the periods (and modes) as well as of the photometry: ‘SOS’ means that the periods and the photometry come from $Gaia$ DR3 using the specific pipeline for the DCEP variables (see Ripepi et al., 2023); ‘DR3’ means that the photometry is from the general DR3 source (see Gaia Collaboration et al., 2023), while ‘P21’ means that the periods are from Pietrukowicz et al. (2021). Note that the distances for the four stars most likely classified as type II Cepheids are probably wrong (see Appendix A). A portion of the Table is shown here for guidance regarding its form and content. The machine-readable version of the full table will be published at the Centre de Données astronomiques de Strasbourg (CDS, https://cds.u-strasbg.fr/).
@@
In recent years, a still unresolved debate has taken place regarding the well-known discrepancy between the values of $H_{0}$ obtained by the SH0ES [^2] Supernovae, HO, for the Equation of State of Dark energy project through the cosmic distance ladder ($H_{0}=$73.01$\pm$0.99 km s-1 Mpc-1, Riess et al., 2022) and the value computed by the Planck cosmic microwave background (CMB) project based on the flat $\Lambda$ cold dark matter ($\Lambda$CDM) model ($H_{0}=$67.4$\pm$0.5 km s-1 Mpc-1, Planck Collaboration et al., 2020). If no solution is found for this intricate $H_{0}$ tension, this could hint at the need for a revision of the $\Lambda$CDM model. It is therefore critical to evaluate the discrepancy through the analysis of the residual systematics (see e.g. Dainotti et al., 2021; Freedman, 2021; Riess et al., 2021, and reference therein).
@@
Metallicity may play an important role as one of the possible residual sources of uncertainty in the cosmic distance ladder, influencing the calibration of the DCEP $PL$ relations. Since the shape and width of the DCEP instability strip are predicted to be affected by metallicity variations (e.g. Caputo et al., 2000), a direct consequence should be seen when estimating the $PL$ relations coefficients (Marconi et al., 2005, 2010; De Somma et al., 2022, and references therein). To avoid systematic effects in the calibration of the extragalactic distance scale (e.g. Romaniello et al., 2008; Bono et al., 2010; Riess et al., 2016), one has to also consider the metallicity dependence of the $PL$ relations and the reddening-free Wesenheit magnitudes (Madore, 1982) in those cases in which the metallicity effect might be small, such as when dealing with the near-infrared regime (NIR, see e.g. Fiorentino et al., 2013; Gieren et al., 2018).
@@
In the past, the lack of accurate independent distances hampered direct empirical evaluations of the metallicity dependence of $PL$ relations using Galactic DCEPs with precise [Fe/H] measurements based on high-resolution (HiRes hereafter) spectroscopy, but this scenario completely changed with the advent of the Gaia mission (Gaia Collaboration et al., 2016). Accurate parallaxes were provided with data release 2 (DR2, Gaia Collaboration et al., 2018) and further improved with the early data release 3 (EDR3, Gaia Collaboration et al., 2021), in addition to the discovery of hundreds of new Galactic DCEPs (Clementini et al., 2019; Ripepi et al., 2019, 2022). A substantial sample can thus be considered when other surveys are taken into consideration, such as the Optical Gravitational Lensing Experiment (OGLE) Galactic Disk survey (Udalski et al., 2018) and the Zwicky Transient Facility (ZTF, Chen et al., 2020), making it possible to improve not only studies related to the cosmic distance scale but also Galactic studies (e.g. Lemasle et al., 2022; Trentin et al., 2023, and references therein). Another constraint was the limited availability of HiRes measurements confined to the solar neighbourhood. This restriction resulted in a narrow range of [Fe/H] values centred on solar or slightly supersolar values, with a small dispersion of 0.2-0.3 dex (e.g. Genovali et al., 2014; Luck, 2018; Groenewegen, 2018; Ripepi et al., 2019). Consequently, this limitation hindered the ability to generate results with statistical significance.
@@
To overcome these problems, we started a project named C-MetaLL [^3] https://sites.google.com/inaf.it/c-metall/home (Cepheid - Metallicity in the Leavitt Law, see Ripepi et al., 2021, for a full description), with the primary goal of measuring the chemical abundance of a sample of at least 300 Galactic DCEPs through HiRes spectroscopy, as well as providing homogeneous multi-band time-series photometry. The main aim was to enlarge the iron abundance range towards the metal-poor regime – that is, [Fe/H]<absent<-< -0.4 dex – in which only a few Galactic stars had abundance measurements in the literature. In the first two papers of the series (Ripepi et al., 2021; Trentin et al., 2023) (R21 and T23 from now on, respectively), we published accurate abundances for more than 25 chemical species for a total of 114 DCEPs, while the first results for the photometric part were presented in Bhardwaj et al. (2024) . In particular, in T23, we obtained measurements for 43 objects with [Fe/H]<absent<-< -0.4 dex, reaching abundances as low as --1.1 dex. This paper represents the sixth manuscript in the series and a direct follow-up of T23, presenting the spectroscopic analysis of 180 new stars based on a total of 331 HiRes spectra and the study of the galactic gradient for 29 chemical elements.
@@
While the main scope of this work is to study the metallicity dependence of the DCEP $PL$ relations in the context of the extragalactic distance scale, the sample of DCEPs presented in this work, in conjunction with our previous results, nevertheless amounts to a total of 294 DCEPs with a homogeneous derivation of stellar parameters and abundances. In addition, as we shall see later in this article, our sample is evenly distributed in metallicity, spanning a range from +0.5 dex to --1 dex in [Fe/H] and 5 to 20 kpc in Galactocentric radius. The aforementioned characteristics make our sample an appropriate testing ground for the study of the Galactic disc abundance gradient, in comparison with literature works that in general use heterogeneous spectroscopic samples or datasets heavily unbalanced towards the solar vicinity (Luck, 2018; Lemasle et al., 2018; Minniti et al., 2020; Ripepi et al., 2022; da Silva et al., 2023; Matsunaga et al., 2023).
@@
The paper is organised as follows: in Sect. 2, we describe the sample of DCEPs and their properties; in Sect. 3, we describe the analysis technique; in Sect. 4 and Sect. 5, we describe and discuss our results; and in Sect. 6, we outline our conclusions.
@@ 2.1 Sample
The sample of DCEPs presented in this work was selected from the DCEP catalogue published in $Gaia$ data release 3 (DR3 Ripepi et al., 2023). We chose the pulsators with the intention of maximising the number of objects with low metallicity and long periods. This is because the main scope of the C-MetaLL survey is to derive accurate $PL$ and $PW$ relations for Galactic DCEPs. To this aim, we have had to enlarge the range covered by the independent variables in the quoted relations as much as possible; namely, periods and metallicities. Therefore, we have had to observe both short- and long-period DCEPs as well as metal-rich and metal-poor pulsators. While the periods are known, the metallicity can only be guessed based on the position of the stars in the disc, which, as was mentioned in the previous section, is well known to exhibit a metallicity gradient from the centre to the anticentre of the Galaxy. Thus, our observations aimed to maximise the most distant objects in the anticentre direction, possibly those with longer periods, which are rarer (since their evolutionary times are faster than those of shorter-period DCEPs). The list of the 180 targets is shown in Table 1, in which we list the main characteristics of the stars, including their periods, modes of pulsation, and distances. These last quantities were calculated based on the $PWZ$ relation published by Ripepi et al. (2022) with the procedure outlined in detail in Sect. 2.2 of Gaia Collaboration et al. (2023). In brief, we used $Gaia$ photometry (see Table 1) to calculate the apparent Wesenheit magnitude $W=G-1.90\times(G_{BP}-G_{RP}$) (Ripepi et al., 2019). Then, we inserted the periods and [Fe/H] values (measured in this paper) in the quoted $PWZ$ to obtain the absolute $W$ magnitude. The distance calculation was then straightforward, using the definition of distance modulus. The conversion in Galactocentric radii $R_{GC}$ was obtained in the usual way, adopting a distance of the Sun from the Galactic centre, $R_{\odot}=8277\pm 9$(stat) ±$\pm 30$(sys) pc (Gravity Collaboration et al., 2022) (for full details, we refer the reader to Gaia Collaboration et al. (2023)).
The distribution of the targets along the Milky Way (MW) disc is displayed in Fig. 1. The distribution of the programme stars ranges from about 5 to 18 kpc in Galactocentric radii ($R_{GC}$) and covers all four Galactic quadrants, with a significant concentration in the third. In the figure, the stars are colour-coded according to their periods. The period distribution shows short-period DCEPs at larger Galactic radii, especially in the third quadrant. This is not a coincidence, but because there exists a period-age relation for DCEPs (see e.g. Bono et al., 2005) so that shorter periods indicate larger ages. It has been found that at larger Galactocentric radii the DCEPs are on average older (e.g. Skowron et al., 2019; De Somma et al., 2020, 2021). This explains the lack of long-period DCEPs at large radii in our sample. An additional feature displayed in Fig. 1 is the presence of three stars with polar angles between 135o and 225o; namely, Gaia DR2 4087335043492541696, V532 Sco, and ASAS-SN J165340.10-332041.7. Given the peculiar position along the Galactic disc, these pulsators may be type II Cepheids. An in-depth analysis of these objects is provided in Appendix A. In addition, a fourth star, OGLE CEP-GD-0069, which was originally classified as DCEP in Udalski et al. (2018), turned out to be a BL Her variable according to Ripepi et al. (2023). This star is also discussed in the appendix. Given their uncertain nature, these four stars have been excluded from the following analysis.
Figure 1: Map along the Galactic disc of the programme stars. The points are colour-coded according to the pulsator’s period. The position of the sun is shown with a yellow-black circle. The four stars probably belonging to the type II Cepheid class are shown with black stars (see text).
@@
Star HJD Phase Texp S/N $T_{eff}$ $\log g$ $\xi$ $V_{broad}$ $RV$ Source
days s K dex km s-1 km s-1 km s-1
AA Gem 59269.76384 0.920 750 102 5656 $\pm$ 68 0.92 $\pm$ 0.05 3.0 $\pm$ 0.2 11.0 $\pm$ 1.0 0.9 $\pm$ 0.3 ESPaDOnS
AA Gem 59630.79632 0.833 410 92 5877 $\pm$ 66 1.22 $\pm$ 0.05 3.3 $\pm$ 0.2 12.0 $\pm$ 1.0 0.8 $\pm$ 0.2 ESPaDOnS
AA Gem 59838.14990 0.162 500 80 4887 $\pm$ 68 0.1 $\pm$ 0.05 2.8 $\pm$ 0.2 9.0 $\pm$ 1.0 24.6 $\pm$ 0.2 ESPaDOnS
AD Pup 59269.81690 0.369 750 78 5128 $\pm$ 100 0.38 $\pm$ 0.05 2.7 $\pm$ 0.2 6.0 $\pm$ 1.0 74.0 $\pm$ 0.3 ESPaDOnS
AD Pup 59630.82225 0.918 410 77 5983 $\pm$ 83 0.58 $\pm$ 0.05 3.4 $\pm$ 0.3 10.0 $\pm$ 1.0 68.3 $\pm$ 0.2 ESPaDOnS
 [^4] The different columns report: the name of the star, the heliocentric Julian day on which the spectrum has been collected, the phase, exposure time, S/N per pixel, effective temperature, logarithm of gravity, microturbulent velocity, broadening velocity, heliocentric radial velocity, and the instrument used to collect the HiRes spectroscopy. The phases were calculated adopting periods and epochs of maximum light from the $Gaia$ DR3 catalogue (Ripepi et al., 2023), except for seven stars for which we report details on the phase determination in Appendix B. A portion is shown here for guidance regarding its form and content. The machine-readable version of the full table will be published at the Centre de Données astronomiques de Strasbourg (CDS, https://cds.u-strasbg.fr/).
@@
 [^5] The table is divided into three parts. The first and last columns report the name of the star and the source, respectively (see Table 2 and 1). The other columns report the estimated abundances (and relative errors) in solar terms for the chemical species analysed in this work. The machine-readable version of the full table will be published at the Centre de Données astronomiques de Strasbourg (CDS, https://cds.u-strasbg.fr/).